Let fn(x)= nx/(1+nx) is not uniformly convergent on [0,1]

1
Expert's answer
2022-03-14T17:00:52-0400

Solution


"% MathType!Translator!2!1!LaTeX.tdl!LaTeX 2.09 and later!\n% MathType!MTEF!2!1!+-\n% feaagGart1ev2aaatCvAUfeBSjuyZL2yd9gzLbvyNv2CaerbuLwBLn\n% hiov2DGi1BTfMBaeXatLxBI9gBaerbd9wDYLwzYbItLDharqqtubsr\n% 4rNCHbGeaGqiVu0Je9sqqrpepC0xbbL8F4HqGqFfpeea0xe9Lq-Jc9\n% vqaqpepm0xbba9pwe9Q8fs0-yqaqpepae9pg0FirpepeKkFr0xfr-x\n% fr-xb9adbaqaaeGaciGaaiaabeqaamaabaabaaGcbaGaamOzamaaBa\n% aaleaacaWGUbaabeaakmaabmaabaGaamiEaaGaayjkaiaawMcaaiab\n% g2da9maalaaabaGaamOBaiaadIhaaeaacaaIXaGaey4kaSIaamOBai\n% aadIhaaaGaaCzcaiaaxMaadaWadaqaaiaaicdacaGGSaGaaGymaaGa\n% ay" "{f_n}\left( x \right) = \frac{{nx}}{{1 + nx}}" for the interval "[0, 1]"


For "x=0" ,


"\mathop {\lim }\limits_{n \to \infty } {f_n}\left( x \right) = \mathop {\lim }\limits_{n \to \infty } \frac{{nx}}{{1 + nx}}\\
\mathop {\lim }\limits_{n \to \infty } {f_n}\left( 0 \right) = \mathop {\lim }\limits_{n \to \infty } \frac{{n\left( 0 \right)}}{{1 + n\left( 0 \right)}}\\
 = 0"


For "0 < x \le 1\"


"\mathop {\lim }\limits_{n \to \infty } {f_n}\left( x \right) = \mathop {\lim }\limits_{n \to \infty } \frac{{nx}}{{1 + nx}}\\
\mathop {\lim }\limits_{n \to \infty } {f_n}\left( x \right) = \mathop {\lim }\limits_{n \to \infty } \frac{{nx}}{{n\left( {\frac{1}{n} + x} \right)}}\\
 = \mathop {\lim }\limits_{n \to \infty } \frac{x}{{\frac{1}{n} + x}}\\
 = 1"


Since we can see that the two limits are not equal, therefore "{f_n}\left( x \right) = \frac{{nx}}{{1 + nx}}" , is not uniformly convergent over the interval "[0, 1]" .